# Utilizing and extending superconducting circuit toolbox to simulate analog quantum gravity

### v2

@@ Abstract
There has been considerable effort to simulate quantum gravity features in solid state systems, such as analog black holes or wormholes. However, superconducting circuits have so far received only limited attention in this regard. Moreover, for quantum superpositions of spacetime geometries – a highly contentious notion within the quantum gravity community itself – there currently exist no solid state blueprints. We here show how quantum circuit hardware can implement a variety of classical and quantum spacetime geometries on lattices, by both using established circuit elements and introducing new ones. We demonstrate the possibility of a metric sharply changing within a single lattice point, thus entering a regime where the spacetime curvature itself is trans-Planckian, and the Hawking temperature ill-defined. In fact, our approach suggests that stable, thermal event horizons are incompatible with strictly discrete lattice models. We thus propose to directly probe the evaporation of a wormhole by tracking the accumulation of charge and phase quantum fluctuations over short time scales, which are a robust signature even in the presence of a dissipative environment. Moreover, we present a loop-hole for the typical black/white hole ambiguity in lattice simulations: the existence of exceptional points in the dispersion relation allows for the creation of pure black (or white) hole horizons –- at the expense of radically changing the interior wormhole dynamics. Finally, based on multistable Josephson junctions, we introduce the notion of quantum inductors: circuit elements that can be prepared in a superposition of different inductance values. Such inductors realise regions with a quantum superposition of analog spacetime. The entanglement of signals with the quantum spacetime can be probed by a type of delayed-choice experiment.
I Introduction
@@
There remain however a number of open questions. First, so far for all solid state implementations the metric can change only over a finite “healing length”. If the curvature at the horizon is too weak, the resulting small Hawking temperature may thus be below the threshold of the system’s intrinsic temperature [43] – an obstacle that (at least up to now) seems to have only been overcome for cold atom simulators [41]. Relatedly, systems with an apparent event horizon generically do not have a well-defined ground state, such that any coupling to the environment leads to an instability, making it extremely challenging to distinguish intrinsic radiation due to the horizon (which would actually simulate aspects of a universe with nontrivial spacetime metric) and radiation due to environment induced relaxation. In addition, even in the absence of environment, the closed system can in some platforms be spontaneously unstable [38, 39]. Furthermore, lattice realizations (which ultimately concerns all solid state systems) add a number of interesting but challenging facets, such as a natural (though possibly artifical) resolution of the trans-Planckian problem [44, 45, 46, 47], and most notably, the fact that any event horizon has both black and white hole character, since the dispersion relation in a periodic Brillouin zone must necessarily cross zero at least twice [31]. Finally, within the quantum gravity community, there also exists the highly contentious idea of superpositions of spacetime geometries, originating from attempts at quantizing the metric tensor [48, 49]. In particular, recent theoretical proposals suggest experiments that examine quantum gravitational effects caused by a superposition of different mass configurations [50, 51, 52, 53, 54, 55, 56, 57]. To the best of our knowledge, there currently exist no feasible blueprints for solid state systems capable of simulating quantum superpositions of spacetime geometries.
@@
Superconducting circuits are one of the prime candidates for building large-scale quantum hardware [58, 59, 60]. The behaviour of these circuits and their interactions with quantised electromagnetic fields in the microwave range are described by circuit quantum electrodynamics (cQED), a formalism that reduces circuits to lumped elements whose phase and charge are canonically conjugate [61, 62, 63, 64, 65, 66, 67, 68, 69]. The toolbox of superconducting circuits includes nonlinear elements such as Josephson junctions and linear elements such as capacitors, inductors, and gyrators – a nonreciprocal multi-port element [70, 71, 72, 73, 74, 75]. The same nonreciprocal behaviour also occurs due to topological transitions in the transport degrees of freedom of multiterminal junctions [76, 77, 78, 79, 80, 81]. Over the course of decades, it has been shown that networks of superconducting circuits can give rise to a large number of physical phenomena and imitate various quantum field theories from other domains [82, 83, 84, 85, 86, 87, 88, 89, 90, 34, 91]. However, the exploration of analog gravity phenomena has so far been limited to Hawking-like radiation in solitons of Josephson junction arrays [32, 33, 92, 93, 94, 95, 34, 35] which leave only little control over the shape of the spacetime geometry, or recreation of a limited class of curved spacetimes by means of flux control of SQUID arrays [36, 37].
@@
In this work, we demonstrate that networks of Josephson junctions and gyrators can be used as an engineering tool for simulations of quantum-gravity phenomena with unprecedented tunability capabilities, and generate surprising insights specific to lattice systems. The basic idea is that arrays of these elements result in a (nearly) massless scalar field theory describing a quantum field – the superconducting phases of the charge islands – moving within a spatially varying analog spacetime metric. The latter essentially encompasses the local dispersion relation, i.e., the velocities with which right and left signals move with respect to the laboratory frame.
@@
To create an apparent horizon we need a boundary between the "normal region" (region where the signals can move in both directions) and a wormhole region where the signals only move in one direction, i.e. a region with overtilted dispersion relation. As it turns out such a region can be created with negative inductors, which we propose achieve by adapting recent insights in flux quench of regular Josephson junctions [96, 69]. Tunable $0-\pi$ junctions [97, 98, 99] may offer an alternative pathway towards negative inductances.
[…]
Moreover, we find an instructive loop-hole to the aforementioned ambiguity of black- versus white hole event horizons in lattice systems. Namely, the inductive coupling can go either via nearest or next-to-nearest neighbour nodes. For the former, the dispersion relation inside the wormhole region exhibits an exceptional point, such that the eigenspectrum no longer crosses zero energy twice, but instead takes a detour on the complex plane to satisfy periodicity. We thus realize lattice versions of event horizons which are definitely either black- or white hole, but not both. However, this feature radically changes the dynamics in the interior of the wormhole: instead of an evaporation from the horizons outwards the entire wormhole interior evaporates everywhere immediately after the quench.
Finally, we go beyond the simulation of classical apparent horizons, and demonstrate that the cQED platform suggested here allows for the implementation of superpositions of spacetime geometries. This idea draws from the recent development of multistable Josephson junctions [100, 101, 102], i.e., junctions whose energy-phase relationship has more than one local minimum. Specifically, the engineering of multiple minimas with different local curvature creates a type of quantum inductor – an element that can be prepared in a superposition of different inductance values. When creating a chain of maximally entangled quantum inductors, we propose to create a superposition of different spacetime geometries. A classical signal traversing this region propagates in a superposition of different velocities with respect to the lab frame, and thus entangles with the quantum spacetime. The entanglement can be probed by a delayed-choice type of experiment, where a Hadamard gate is applied on the quantum inductor array after traversal of the quantum spacetime region, leading to a specific interference pattern.
In an extended outlook, we show that in particular the idea of quantum superpositions of spacetime geometries harbors a great potential for follow-up studies, where one could endow the spacetime itself with a nonzero Hamiltonian, thus working towards much more complex simulations of analog quantum gravity, where both the field moving in curved analog spacetime and the spacetime itself have a quantum coherent and strongly correlated dynamics. Moreover, while we mostly consider long chains in the main part of this work, as they provide a clean interpretation of the effects in the context of curved spacetime, we are aware that they may be challenging to realize experimentally. In the outlook, we therefore also propose experiments on only few lattice points as proofs of principle, which already contain much of the pertinent phenomenology present in long chains.
@@
This paper is organized as follows, section II is a brief review of scalar field theories on curved spacetime, and a summary of our main accomplishments of this work. In section III we introduce and study the two circuits that will host the apparent horizons that we wish to investigate, we especially focus on the different effect that negative inductances have on the dispersion relations of these two circuits anticipating the differences between the horizons in them. Section IV is a brief detour, where we expand upon the idea of using flux quench on a Josephson junction to get a negative inductance, and in section V we study and characterize the horizons in the two aforementioned circuits, by looking at the time evolution of quantum fluctuations of the conjugate charge and the phase difference. In section VI we hypothesize on the long time fate of the analog horizons and argue that the effect of this intrinsic evaporation is easily distinguishable from the effect of coupling to the environment. Finally, in section VII we present our proposal for observing quantum superpositions of two spacetime geometries using multi-stable Josephson junctions.
II Curved space time metrics and the circuit simulator
@@
II.1 Brief review of field theories with curved spacetime and stability considerations
@@
Within this brief review, questions regarding the stability of the considered systems will emerge. To this end, we begin by providing a number of general statements regarding bosonic Hamiltonians (as we focus on scalar bosonic quantum fields). The quantum systems we consider are all described by a Hermitian Hamiltonian. In accordance with Refs. [38, 39], we point out that hermiticity does not guarantee stability of the system, due to the special symplectic structure of the Bogoliubov transformation. In particular, two cases need to be distinguished. It may happen that certain eigenvalues are negative, such that the system has no well-defined many-body ground state. Such systems however still evolve in a stable fashion, unless they are coupled to an environment, which will in general lead to a collapse of the system, as negative energy bosonic states can now be occupied by extracting energy from the bath. In the second case, some eigenenergies may become complex (again, this is consistent with Hermitian Hamiltonians, see Refs. [38, 39]). Such systems are spontaneously unstable, as they do not require a coupling to a bath to collapse.
In continuous field theories, the action of a field $\phi$ in a ${d+1}$-dimensional spacetime is generally given as
[…]
@@
For simplicity, let us consider ${1+1}$ dimensions (note that our simulator recipe works also for more spatial dimensions, as we point out in a moment), and focus on metrics that depend only on the position coordinate, i.e., metrics that are stationary in the laboratory frame. Here we write the action explicitly in a matrix form as
[…]
where $u$ and $v$ are two independent functions of $x$. Note that the matrix representing the spacetime metric has determinant $-1$, such that the $\sqrt{-g}$ prefactor is here irrelevant. For actions of this form, it is always possible to make a coordinate transformation [43]
[…]
@@
In these coordinates, the classical Euler-Lagrange equation is simply ${\partial_{y}\partial_{z}\phi=0}$, whose solutions are arbitrary superpositions of functions that depend either only on $y$, or only on $z$. Transforming back to $(t,x)$-coordinates via Eqs. (3) and (4), these two separate solutions correspond to waves propagating locally either with velocity $u+v$ or $u-v$. For constant $u,v$ (and $u>0$), the theory is readily quantized, yielding the Hamiltonian
[…]
@@
where for bosonic fields, $[a_{k},a_{k^{\prime}}^{\dagger}]=\delta(k-k^{\prime})$. The dispersion relation $\omega_{k}=u|k|+vk$ readily provides the group velocity of quantum excitations, such that the classical $y$ and $z$ type solutions here correspond to $k>0$ and $k<0$, respectively. For $|v|<u$, the system describes a regular analog light cone with $H$ having only positive many-body eigenvalues, where $v$ corresponds to a finite tilt of the dispersion relation.
@@
For $|v|>u$ the dispersion relation is overtilted, such that one of the branches (either $k<0$ or $k>0$) has now negative energies. Negative energies are meaningful in the following sense. As already pointed out above, the Hamiltonian $H$ here has in principle no more well-defined ground state – indicating a serious instability. However, the eigenvalues remain real, such that the closed quantum system still evolves in a stable fashion. In particular, with a Galilei transformation  [^1] Galilei transformations are of course allowed for the here considered simulator systems which operate at velocities far below the speed of light. by, e.g., $-v$, we can undo the tilt (the dispersion relation goes to $u|k|$) and thus return to a regular Hamiltonian with a well-defined ground state. Consequently, for the closed quantum system, there is nothing special about whether or not a particular branch has negative eigenenergies. Matters are markedly different, once we consider open quantum systems. Here, the Galilei transformation applies to both the system Hamiltonian, as well as to the interaction with the environment. Hence, one cannot change a system from stable to unstable by merely moving with respect to it – thus avoiding what would otherwise be a serious conundrum. However, the inverse may well happen: if we fine tune the parameters of the system such that it implements an overtilted dispersion relation in the laboratory rest frame, the coupling to the environment may indeed be such that the system collapses on very fast time scales (usually given by the typical relaxation rates of the considered quantum hardware).
@@
Returning to spatially varying spacetime metrics, apparent event horizons emerge in this field theory when at a given point $v$ crosses from $|v|<|u|$ to $|v|>|u|$. Consequently, the system is separated at the horizon into two halves, one with a well-defined ground state (regular or no tilt in the dispersion relation) and one with no ground state (overtilted). In principle, one could now expect already the closed quantum system to be unstable as the two halves could exchange energy (which would thus lead to complex eigenvalues). But as is well known [43], this does not happen for this particular system. Even in the presence of such a horizon, the system is still described by regular, real eigenvalues, such that at least for the closed system, the evolution is still stable (though in general not in a well-defined ground state). The reason for this is that on both sides of the horizon, the above coordinate transformation to $(y,z)$ still applies, such that for these coordinates, there still exist two well-defined branches of the wave solutions moving at different speeds relative to the lab frame. At the horizon, the transition from $|v|<|u|$ to $|v|>|u|$ results in one – but only one – of the two coordinate transformations (either $y$ or $z$, depending on the sign of $v$) to be singular. Consequently, the branch without singularity simply moves through the event horizon as if it wasn’t there. The other branch (with the singularity) is causally disconnected at the horizon, because wave solutions come to a stand-still. Hence the left and right hand side of this branch do not couple.
@@
Since the system is however not in a ground state, meaningful quantum measurements are best defined by anchoring all observables to a basis with well-defined ground state. This is commonly done [43] by imagining the system to be prepared (in the distant past) in the ground state of a spacetime metric without horizon (e.g., perfectly flat spacetime). When now defining measurements with respect to the new eigenbasis, Hawking radiation emerges. A frequency decomposition of this radiation reveals that it is of thermal nature, given by the curvature at the horizon – e.g., for constant $u$, the temperature is thus proportional to the energy scale $\partial_{x}v$, evaluated at the horizon.
@@
II.2 Simulation with circuit networks

### v4

# On-demand analog space-time in superconducting networks: grey holes, dynamical instability and exceptional points

@@ Abstract
There has been considerable effort to mimic analog black holes and wormholes in solid state systems. Lattice realizations in particular present specific challenges. One of those is that event horizons in general have both white and black hole (grey hole) character, a feature guaranteed by the Nielsen-Ninomiya theorem. We here explore and extend the capability of superconducting circuit hardware to implement on-demand spacetime geometries on lattices, combining nonreciprocity of gyrators with the nonlinearity of Josephson junctions. We demonstrate the possibility of the metric sharply changing within a single lattice point, thus entering a regime where the modulation of system parameters is “trans-Planckian”, and the Hawking temperature ill-defined. Instead of regular Hawking radiation, we find an instability in the form of an exponential burst of charge and phase quantum fluctuations over short time scales – a robust signature even in the presence of an environment. Moreover, we present a loop-hole for the typical black/white hole ambiguity in lattice simulations: exceptional points in the dispersion relation allow for the creation of pure black (or white) hole horizons, at the expense of a radical change in the dynamics of the wormhole interior.
Copyright attribution to authors.
This work is a submission to SciPost Physics.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
1 Introduction
@@
There remain however a number of open questions. First, so far for solid state implementations the metric usually changes only over a finite “healing length”. If the surface gravity (for analog black holes surface gravity corresponds to the rate of change of the group velocity with which a signal will move in the given metric [41]) is too weak, the resulting small Hawking temperature may thus be below the threshold of the system’s intrinsic temperature [43] – an obstacle that (at least up to now) seems to have only been overcome for cold atom simulators [41]. Relatedly, systems with an apparent event horizon generically do not have a well-defined ground state, such that any coupling to the environment leads to an instability, making it extremely challenging to distinguish intrinsic radiation due to the horizon (which would actually simulate aspects of a universe with nontrivial spacetime metric) and radiation due to environment induced relaxation. In addition, even in the absence of environment, the closed system can in some platforms be spontaneously unstable [38, 39]. Furthermore, lattice realizations (which ultimately concerns all solid state systems) add a number of interesting but challenging facets, such as a natural (though possibly artifical) resolution of the trans-Planckian problem [44, 45, 46, 47], and most notably, the fact that any event horizon has both black and white hole character (also referred to as grey hole), since the dispersion relation in a periodic Brillouin zone must necessarily cross zero at least twice [31], a simple consequence of the Nielsen-Ninomiya theorem [48].
@@
Superconducting circuits are one of the prime candidates for building large-scale quantum hardware [49, 50, 51]. The behaviour of these circuits and their interactions with quantised electromagnetic fields in the microwave range are described by circuit quantum electrodynamics (cQED), a formalism that reduces circuits to lumped elements whose phase and charge are canonically conjugate [52, 53, 54, 55, 56, 57, 58, 59, 60]. The toolbox of superconducting circuits includes nonlinear elements such as Josephson junctions and linear elements such as capacitors and inductors. Our work will also feature nonreciprocal multi-port circuit elements. Classical nonreciprocal elements called gyrators (or their close cousins, circulators), are already widely in use in circuit engineering and signal processing [61, 62, 63, 64], whose consistent circuit theoretical description goes back to the work by Tellegen [65]. These are however generally large clunky objects operating in a finite frequency window. Various recent works strive towards a realisation and a consistent description of quantum mechanical gyrators [66, 67, 68, 69]. Moreover, it has recently been realised [69, 70, 71, 72] that the same nonreciprocal behaviour also occurs due to topological transitions in the transport degrees of freedom of multiterminal junctions  [73, 74, 75, 76, 77, 78, 79, 80, 81], whose size can be on the meso scale and the gyration behavior emerges in the dc limit.
Over the course of decades, it has been shown that networks of superconducting circuits can give rise to a large number of physical phenomena and imitate various quantum field theories from other domains [82, 83, 84, 85, 86, 87, 88, 89, 90, 34, 91, 92, 93, 94]. However, the exploration of analog gravity phenomena has so far been limited to Hawking-like radiation in Josephson junction arrays [32, 33, 95, 96, 97, 98, 34, 35], which, due to the nature of their proposal leave only little control over the shape of the spacetime geometry, or recreation of a limited class of curved spacetimes by means of flux control of SQUID arrays [36, 37].
@@
In this work, we demonstrate that networks of Josephson junctions and gyrators can be used as an engineering tool for simulations of quantum-gravity phenomena with unprecedented tunability capabilities, and generate surprising insights specific to lattice systems. The basic idea is that arrays of these elements result in a (nearly) massless scalar field theory describing a quantum field – the superconducting phases of the charge islands – propagating across a lattice with spatially varying analog spacetime metric. Contrary to effective modulations of spacetime (metric) by the excitation of solitons [98] in superconducting circuits, here a nontrivial spacetime and event horizons emerge via direct parametric control of the gyration and induction values of the circuit elements within the network. The local analog spacetime metric essentially encompasses the local dispersion relation, i.e., the velocities with which right and left signals move with respect to the laboratory frame.
@@
To create an apparent horizon we need a boundary between the "normal region" (region where the signals can move in both directions) and a wormhole region where the signals only move in one direction, i.e., a region with overtilted dispersion relation. As it turns out such a region can be created with negative inductors, which we propose to achieve by adapting recent insights in flux quench of regular Josephson junctions [99, 60]. Tunable $0-\pi$ junctions [100, 101, 102] may offer an alternative pathway towards negative inductances.
Already for such classical metric profiles, we make a number of important observations specific to lattice simulations. The exceptional tunability of the individual energy scales in the lattice allows to change the analog metric over only a few lattice sites, which would yield for typical charging and Josephson energies Hawking temperatures in the $100\text{mK}$ up to $1K$ range, thus comfortably exceeding usual cryogenic temperatures. But as we show, this comes at the price of a fundamental impossibility to create stable event horizons for strictly discrete lattice systems. Instead, the system becomes unstable right after the quench, resulting in an immediate evaporation starting out from the event horizon. We show that the instability leads to an accumulation of charge and phase quantum fluctuations, which we expect to be a highly robust signature with respect to environment-induced dissipation, since the latter reduces quantum fluctuations (instead of increasing them). Moreover, we find an instructive loop-hole to the aforementioned ambiguity of black- versus white hole event horizons in lattice systems. Namely, the inductive coupling can go either via nearest or next-to-nearest neighbour nodes. For the former, the dispersion relation inside the wormhole region exhibits an exceptional point (EP), such that the eigenspectrum no longer crosses zero energy twice, but instead takes a detour on the complex plane to satisfy periodicity. We thus realize lattice versions of event horizons which are definitely either black- or white hole, but not both. However, this feature radically changes the dynamics in the interior of the wormhole: instead of an evaporation from the horizons outwards the entire wormhole interior evaporates everywhere immediately after the quench. Our work thus connects to the currently highly active research field dealing with EPs, generic topological defects in non-Hermitian matrices [103]. EPs have been studied in various areas of physics such as optics [104, 105], acoustics [106] and quantum mechanics [107, 108, 109]. They also play an important role in condensed matter systems described by non-Hermitian Hamiltonians [110, 111, 112], specifically as a basis for their topological properties [113, 114, 115, 116, 117]. In our work the non-Hermitian matrix (and EPs) arise due to bosonic Bogoliubov transformations. We will go into a little more detail about complex eigenvalues as we encounter them, but in short this is due to the fact that the Bogoliubov transformations for bosons are symplectic, unlike fermions where the transformations are unitary. Furthermore, while we consider long chains in the main part of this work, as they provide a clean interpretation of the effects in the context of curved spacetime, we are aware that they may be challenging to realize experimentally. Therefore in the outlook, we propose experiments on only few lattice points as proofs of principle, which already contain much of the pertinent phenomenology present in long chains.
The prediction of both astrophysical and analog event horizons is predicated on the semi-classical approximation, i.e. quantum field theory on fixed background spacetime. Our work here provides a first step towards investigating the backaction of quantum fields on spacetime, as should be clear from our discussion in Sec.6, where we qualitatively consider the long-term effect of decaying event horizons in our system. However, this point needs further investigation in future with some concrete calculations, possibly with a simplified model of the effect of evolving quantum fields on spacetime. We have planned a manuscript in the future with one possible realisation of this idea in superconducting circuits using multi-stable JJs [118, 119, 120].
@@
This paper is organized as follows. Section 2 is a brief review of scalar field theories on curved spacetime, and a summary of our main accomplishments of this work. In Sec. 3 we introduce and study the two circuits that will host the apparent horizons that we wish to investigate, we especially focus on the different effect that negative inductances have on the dispersion relations of these two circuits anticipating the differences between the horizons in them. Section 4 is a brief detour, where we expand upon the idea of using flux quench on a Josephson junction to get a negative inductance, and in Sec. 5 we study and characterize the horizons in the two aforementioned circuits, by looking at the time evolution of quantum fluctuations of the conjugate charge and the phase difference. In Sec. 6 we hypothesize on the long time fate of the analog horizons and argue that the effect of this intrinsic evaporation is easily distinguishable from the effect of coupling to the environment.
2 Curved space time metrics and the circuit simulator
@@
2.1 Brief review of field theories with curved spacetime and stability considerations
@@
Within this brief review, questions regarding the stability of the considered systems will emerge. To this end, we begin by providing a number of general statements regarding bosonic Hamiltonians (as we focus on scalar bosonic quantum fields). The quantum systems we consider are all described by a Hermitian Hamiltonian. In accordance with Refs. [38, 39, 121, 122], we point out that hermiticity does not guarantee stability of the system, due to the special symplectic structure of the Bogoliubov transformation. In particular, two cases need to be distinguished. It may happen that certain eigenvalues are negative, such that the system has no well-defined many-body ground state. Such systems however still evolve in a stable fashion, unless they are coupled to an environment, which will in general lead to a collapse of the system, as negative energy bosonic states can now be occupied by extracting energy from the bath. In the second case, some eigenenergies may become complex (again, this is consistent with Hermitian Hamiltonians, see Refs. [38, 39]). Such systems are spontaneously unstable, as they do not require a coupling to a bath to collapse.
In continuous field theories, the action of a field $\phi$ in a d+1{d+1}-dimensional spacetime is generally given as
$S=\frac{1}{2}\int dtd^{d}x\sqrt{-g}\,g^{\mu\nu}\partial_{\mu}\phi\,\partial_{\nu}\phi\ ,$ (1)
@@
For simplicity, let us consider 1+1{1+1} dimensions (note that our simulator recipe works also for more spatial dimensions, as we point out in a moment), and focus on metrics that depend only on the position coordinate, i.e., metrics that are stationary in the laboratory frame. Here we write the action explicitly in a matrix form as
$S=\frac{1}{2}\int dtdx\left(\begin{array}[]{cc}\partial_{t}\phi&\partial_{x}\phi\end{array}\right)\left(\begin{array}[]{cc}\frac{1}{u}&\frac{v}{u}\\ \frac{v}{u}&\frac{v^{2}-u^{2}}{u}\end{array}\right)\left(\begin{array}[]{c}\partial_{t}\phi\\ \partial_{x}\phi\end{array}\right)\ ,$ (2)
where $u$ and $v$ are two independent functions of $x$. Note that the matrix representing the spacetime metric has determinant 1-1, such that the $\sqrt{-g}$ prefactor is here irrelevant. For actions of this form, it is always possible to make a coordinate transformation [43]
$y$ $=t-\int^{x}dx^{\prime}\frac{1}{u\left(x^{\prime}\right)+v\left(x^{\prime}\right)}$ (3)
$z$ $=t+\int^{x}dx^{\prime}\frac{1}{u\left(x^{\prime}\right)-v\left(x^{\prime}\right)}\ ,$ (4)
@@
In these coordinates, the classical Euler-Lagrange equation is simply $\partial_{y}\partial_{z}\phi=0$}, whose solutions are arbitrary superpositions of functions that depend either only on $y$, or only on $z$. Transforming back to $(t,x)$-coordinates via Eqs. (3) and (4), these two separate solutions correspond to waves propagating locally either with velocity $u+v$ or uvu-v. For constant $u,v$ (and $u>0$), the theory is readily quantized, yielding the Hamiltonian
$H=\int dk\omega_{k}a_{k}^{\dagger}a_{k}\ ,$ (6)
@@
where for bosonic fields, $a_{k},a_{k^{\prime}}^{\dagger}]=\delta(k-k^{\prime})$. The dispersion relation $\omega_{k}=u|k|+vk$ readily provides the group velocity of quantum excitations, such that the classical $y$ and $z$ type solutions here correspond to $k>0$ and $k<0$, respectively. For $|v|<u$, the system describes a regular analog light cone with $H$ having only positive many-body eigenvalues, where $v$ corresponds to a finite tilt of the dispersion relation.
@@
For $|v|>u$ the dispersion relation is overtilted, such that one of the branches (either $k<0$ or $k>0$) has now negative energies. Negative energies are meaningful in the following sense. As already pointed out above, the Hamiltonian $H$ here has in principle no more well-defined ground state – indicating a serious instability. However, the eigenvalues remain real, such that the closed quantum system still evolves in a stable fashion. In particular, with a Galilei transformation  [^1] Galilei transformations are of course allowed for the here considered simulator systems which operate at velocities far below the speed of light. by, e.g., v-v, we can undo the tilt (the dispersion relation goes to $u|k|$) and thus return to a regular Hamiltonian with a well-defined ground state. Consequently, for the closed quantum system, there is nothing special about whether or not a particular branch has negative eigenenergies. Matters are markedly different, once we consider open quantum systems. Let us stress though that the Galilei transformation applies to both the system Hamiltonian, as well as to the interaction with the environment. Hence, one cannot change a system from stable to unstable by merely moving with respect to it – thus avoiding what would otherwise be a serious conundrum. However, the inverse may well happen: if we fine tune the parameters of the system such that it implements an overtilted dispersion relation in the laboratory rest frame, the coupling to the environment may indeed be such that the system collapses on very fast time scales (usually given by the typical relaxation rates of the considered quantum hardware).
@@
Returning to spatially varying spacetime metrics, apparent event horizons emerge in this field theory when at a given point $v$ crosses from $|v|<|u|$ to $|v|>|u|$, such that there no longer exist a global transformation to a flat spacetime. Consequently, the system is separated at the horizon into two halves, one with a well-defined ground state (regular or no tilt in the dispersion relation) and one with no ground state (overtilted). In principle, one could now expect already the closed quantum system to be unstable as the two halves could exchange energy (which would thus lead to complex eigenvalues). But as is well known [43], this does not happen for this particular system. Even in the presence of such a horizon, the system is still described by regular, real eigenvalues, such that at least for the closed system, the evolution is still stable (though in general not in a well-defined ground state). The reason for this is that on both sides of the horizon, the above coordinate transformation to $(y,z)$ still applies, such that for these coordinates, there still exist two well-defined branches of the wave solutions moving at different speeds relative to the lab frame. At the horizon, the transition from $|v|<|u|$ to $|v|>|u|$ results in one – but only one – of the two coordinate transformations (either $y$ or $z$, depending on the sign of $v$) to be singular. Consequently, the branch without singularity simply moves through the event horizon as if it wasn’t there. The other branch (with the singularity) is causally disconnected at the horizon, because wave solutions come to a stand-still. Hence the left and right hand side of this branch do not couple.
@@
Since the system is however not in a ground state, meaningful quantum measurements are best defined by anchoring all observables to a basis with well-defined ground state. This is commonly done [43] by imagining the system to be prepared (in the distant past) in the ground state of a spacetime metric without horizon (e.g., perfectly flat spacetime). When now defining measurements with respect to the new eigenbasis, Hawking radiation emerges. A frequency decomposition of this radiation reveals that it is of thermal nature, given by the surface gravity – e.g., for constant $u$, the temperature is thus proportional to the energy scale $\partial_{x}v$, evaluated at the horizon.
@@
2.2 Simulation with circuit networks